OpenStudy (anonymous):

write as the power of 5 in simplest form: 5 sqrt(5) 1/ 4 sqrt(5) 25^a+3

OpenStudy (anonymous):

5sqrt(5)=sqrt125 sqrt(5/16) (25^a)*(25^3)

OpenStudy (phi):

\[5\sqrt{5}= 5\cdot 5^{\frac{1}{2}}= 5^{\frac{3}{2}}\] Can you do (1/4) sqrt(5) ? 25 is 5^2 so \[ 25^{(a+3)}= (5^2)^{(a+3)}\] can you simplify the last even more??